OpenStudy (anonymous):

write an equation in point-slope form for the line through the given point with the given slope. (-3, -5); m= -2/5 A)y+= -2/5(x-8) B)y+5= (-2/5)x-3 C)y-5=-2/5(x+3) D) )y-5=-2/5(x-3)

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

OpenStudy (anonymous):

y-y1=m(x-x1) so y-(-5)=(-2/5)(x-(-3))